### v1

# DreamBench++: A Human-Aligned Benchmark for Personalized Image Generation

Abstract
Personalized image generation holds great promise in assisting humans in everyday work and life due to its impressive function in creatively generating personalized content. However, current evaluations either are automated but misalign with humans or require human evaluations that are time-consuming and expensive. In this work, we present DreamBench++, a human-aligned benchmark automated by advanced multimodal GPT models. Specifically, we systematically design the prompts to let GPT be both human-aligned and self-aligned, empowered with task reinforcement. Further, we construct a comprehensive dataset comprising diverse images and prompts. By benchmarking 7 modern generative models, we demonstrate that DreamBench++ results in significantly more human-aligned evaluation, helping boost the community with innovative findings.
Figure 1: Overview of DreamBench++. We collect diverse images and prompts, and utilize GPT-4o for automated evaluation aligned with human preference.
1 Introduction
Driven by the significant advances in large-scale text-to-image (T2I) generative models [85, 82, 8, 83, 69, 88, 116, 14, 27, 21, 22, 7, 49, 24], it is now possible to generate images conditioned on not only arbitrary text prompts but also by given reference images—personalized image generation [86, 28, 55, 115, 53, 29, 6, 17, 48, 15, 111, 98, 109, 67, 44, 105, 66, 103, 16, 99, 126]. In general, to be useful as an artistic creation tool for inspiration or products [113], the following two basic criteria must be fulfilled: i) Prompt following (image & prompt consistency). Generated images must follow the prompt description, which is a requirement shared with vanilla T2I generation [8, 83]. ii) Concept preservation (image & image consistency). For personalized image generation, the concept of the reference image, i.e., the main subject’s semantic details (e.g., facial characters) or high-level abstractions (e.g., overall style), must be preserved in the generated image. For example, a user may want to “imagine his own dog traveling around the world” [86], and the generated dog must be the same as his but traveling.
Figure 2: Qualitative comparison of concept preservation evaluation between DreamBench++ and traditional DINO [13]. DINO often fails to yield human-aligned evaluation while our DreamBench++ succeeds using multimodal GPT models as the evaluator.
To meet the aforementioned requirements, numerous efforts have been devoted. One line of fine-tuning-based works focuses on fine-tuning general T2I models to specialist personalization models by reproducing specific concepts present in training sets [86, 28, 17, 53]. Meanwhile, another line of encoder-based works, instead, achieves concept-preservation by training features adaptation to inject reference image features into a general T2I model [115, 29, 6, 24, 95, 96, 73]. Although remarkable results have been achieved, one question arises: can we comprehensively evaluate these models to figure out which technical route is superior and where to head?
In this work, we aim to answer this question by developing a new benchmark that properly evaluates personalized T2I models driven by the above two requirements. We present DreamBench++, a comprehensive benchmark designed based on the following de-facto principled advantages:
  • 1.
    Human-Aligned  As shown in Fig. 2, traditional metrics like DINO [13] and CLIP [81] often result in significant discrepancies from humans. This is caused by the image similarity measurement nature of DINO and CLIP models, and thus crowd-sourced human evaluation is typically necessary for obtaining a correct quantitative understanding of generated images [54, 52, 112]. Therefore, different from existing works that utilize CLIP and DINO as metrics that may be humanly misaligned, our DreamBench++ demonstrates surprisingly consistent evaluation results aligned with humans. For instance, by evaluating 7 modern models, DreamBench++ achieves 79.64% and 93.18% agreement with human’s evaluation in concept preservation and prompt following capabilities, respectively. Notably, it is +54.1% and +50.7% higher than traditional DINO and CLIP metrics.
  • 2.
    Automated  However, it is non-standardized and expensive to perform high-quality human evaluations, making human evaluations non-trivial to use. To address this challenge, DreamBench++ achieves automated but human-aligned evaluation by using advanced multimodal GPT models as metrics such as GPT-4o [71]. The challenges lie in two aspects: i) prompt design and ii) reasoning procedure for scoring. We systematically standardize the automated GPT evaluation by first designing the evaluation instructions that provide overall task requirements, where language is a general interface for instructing human preference. Inspired by Self-Align [97], we instruct the GPT model to conduct internal thinking that aligns itself for better task and preference understanding. Then, the GPT model provides the summary & planning for the task and scoring criteria, and the final scores are provided with optional chain-of-thought (CoT) [108, 122].
  • 3.
    Diverse  To avoid biased understanding due to the low diversity of evaluation data, DreamBench++ collects a large number of diverse images covering various levels of difficulty categories, from animals and styles that are relatively simple to more challenging human subjects, objects, and non-natural image styles (See Fig. 1). Compared to commonly used DreamBench [86] that only contains 30 subjects and 25 prompts, DreamBench++ scales up the benchmarking data to 150 images and 1,350 prompts, i.e., 5$\times$ and 54$\times$ more than DreamBench. As a result, when evaluating models on this larger dataset, different but more robust conclusions have been obtained that demonstrate an unbiased and more comprehensive evaluation of DreamBench++.
Takeaways  Here, we offer some insightful findings that we obtain by evaluating 7 modern personalized T2I models: i) DINO-based rating puts tremendous weight on resemblance in overall shape or color and neglects detailed visual features, which makes it a sub-optimal metric for personalized image generation evaluation; ii) The main goal of personalized image generation is to find the Pareto optimal solution that balances concept preservation and prompt following. Among these 7 models, DreamBooth [86] is the best in overall performance, preserving highly detailed visual features while keeping close adherence with text prompt; iii) Currently, personalized T2I models perform well on animal and style category. However, they have not performed well in the human category due to sensitivity to facial details and in the object category due to its diversity. Existing works [105, 100, 114, 115, 111] focuses on preserving facial features to address the former problem, but the latter has not been extensively explored.
We are presenting DreamBench++ with open-sourced codes and evaluation standardization to promote innovation within the research community. In addition, we believe our design of the human-aligned & automated evaluation using advanced foundation models is robust and transferrable to other domains and foundation models (e.g., GPT-5 in the future).
2 DreamBench++
We introduce DreamBench++, a human-aligned, automated, and diverse benchmark that evaluates the two capabilities of personalized image generation models. In the following, we describe how we construct DreamBench++ from two aspects: prompts and data.
Figure 3: Overall procedure of prompting GPT-4o for automated evaluation. The evaluation instructions are meta-prompting information written by humans, including task description, scoring criteria, scoring range, and format specification. Then, GPT-4o is prompted with reasoning instructions to perform internal thinking that provides a self-aligned task summary and planning. Finally, all prompts and reasoning outputs are joined with image samples for score outputs.
2.1 Prompting Multimodal GPT for Automated & Human-Aligned Benchmarking
It is challenging to obtain a solid quantitative understanding of generated models, especially when evaluating visual contents like images that rely on human evaluations [54, 52]. Thus, it is critical to achieve automated evaluation by utilizing multimodal GPT models, which are trained particularly in the principle of aligning with human preference [72, 20, 71, 70]. This is evidenced by the recent progress achieved by Wu et al., which demonstrates that GPT-4V [70] can serve as a human-aligned text-to-3D generation evaluator. However, as pointed out by Zhang et al. and Ku et al., multimodal GPT models often fall short in evaluating personalized image generation—often more challenging when distinguishing subtle difference for concept preservation assessment using GPT—still underexplored. To tackle this issue, we detail how we systematically design the prompt of multimodal GPT (GPT-4o [71], by default) for human alignment reinforcement but also improve the reasoning progress that helps the GPT models to be more self-aligned, introduced as follows.
Compare or rate?  There are typically two schemes for quantitatively evaluating generative models in human evaluations: rating and comparison [121, 124]. The rating scheme requires human reviewers to assign an absolute score to each instance, while the comparison scheme asks human reviewers to express a relative preference among different instances. Though effective as the comparison scheme is when humans are involved, we find that there are two critical issues. i) Positional Bias: the scoring results of GPT-4V/GPT-4o is sensitive to the order in which images are presented [70, 104, 107, 121, 110, 124], making it unsuitable for comparison scheme. ii) Quadratic Complexity: As the number of methods increases, the number of essential evaluation runs for numerical rating increases linearly, while the number of comparative assessments increases quadratically. Therefore, direct numerical rating is more efficient and scalable when evaluating multiple methods. Hence, in this work, we adhere to the rating scheme, and we establish a 5-level rating scheme where scores are integers ranging from 0 (very poor) to 4 (excellent).
Figure 4: Dataset construction process of our DreamBench++. We start by obtaining keywords through GPT generation, existing datasets, and human proposals. Next, we collect corresponding images from the internet. These images are then filtered to remove low-quality ones through both model and human assessment. The remaining high-quality images are used as input for GPT-4o to generate text prompts of varying difficulty levels.
Evaluation Instructions  The evaluation instructions serve as the meta-prompting that describes overall tasks, which is shown in Fig. 3. As stated in Section 1, there are two fundamental quality criteria to be evaluated: i) concept preservation and ii) prompt following. For each aspect, we use a similar prompt template that contains ❶ task description, ❷ scoring criteria explanation, ❸ scoring range definition, and ❹ format specification. Only the scoring criteria are tailored for different tasks: for concept preservation evaluation, we prompt GPT to focus on shape, color, texture, and facial features (if applicable), while for prompt following evaluation we requested for focus on relevance, accuracy, completeness and context.
Reasoning Instructions  Given the evaluation instructions, it is crucial to reinforce the alignment with both the human instruction and itself to largely leverage the pretrained knowledge. To this end, we adopt a 2-step evaluation policy as follows: i) Internal Thinking: Inspired by Self-Align [97], we introduce internal thinking to strengthen task understanding and instruction following capabilities. Specifically, we prompt the GPT model by asking if it understands the task or not and let it summarize the task. ii) Summary & Planning: According to the given internal thinking instruction, the GPT will summarize and plan for the evaluation task itself. It can also be viewed as a generalized form of chain-of-thought reasoning [108, 122]. The complete procedure is illustrated in Fig. 3.
2.2 Scaling Up Personalized Image Generation Benchmarking
Pioneering works like DreamBooth [86] and SuTI [17] have successfully set up baseline datasets for the evaluation of personalized image generation, and DreamBench++ follows them to categorize images into three types: ❶ objects, ❷ living subjects, and ❸ styles. However, due to the small-scale nature of DreamBench, it is limited as some methods may converge well on its samples while performing unsatisfactorily on other data. To avoid this possible biased evaluation, we scale up the benchmarking data by increasing both image numbers and diversity.
Figure 5: Data distribution visualization. (a) Images comparison between DreamBench and DreamBench++ using t-SNE [101, 77]. (b) Image and prompt distribution of DreamBench++.
Data Construction from Internet  There are broad images on the Internet, and many datasets are constructed from it [90, 47]. DreamBench++ mainly collects images from Unsplash [3], Rawpixel [2], and Google Image Search [1], along with contributions from individuals with authorized permissions. Each image’s copyright status has been verified for academic suitability. As shown in Fig. 4, we collect and construct high-quality data in DreamBench++ by following 3 steps:
  • Keywords Generation  First, we generate 200 relevant keywords using GPT-4o and join them with the 200 most frequent keywords from Unsplash. After filtering out duplicated keywords, seven human annotators will extend the list to around 300 based on their interests.
  • Internet Images Collection  Given selected keywords, we retrieved corresponding images from Unsplash, Rawpixel, and Google Image Search. To filter out images unsuitable for personalized image generation, SAM [50] is applied to identify subject regions in images and discard those with too small subject areas. Human annotators will then filter out images with noisy backgrounds. Curated images were cropped to centralize the subject, resulting in two images per keyword. Keywords that fail to yield suitable images will be discarded in this process.
  • Prompt Generation  After image collection, 9 text prompts per image were generated using GPT-4o, designed to cover a range of difficulties: 4 prompts for ❶ photorealistic styles, 3 for ❷ non-photorealistic styles, and 2 for ❸ complicated & imaginative contents. To align with established evaluation methods, we use few-shot prompts selected from PartiPrompts [116]. Human calibration ensures that all generated prompts are ethical and without flaws. As a result, the construction process finally yields 150 high-quality images and 1,350 prompts.
Diversity Visualization  When collecting images, we notice that Internet images possess a bias towards photorealistic styles. To diversify, various non-photorealistic styles are enlisted, and human annotators are tasked to gather images for each style, including anime, sketches, traditional Chinese paintings, artworks, and cartoon characters from games. Then, a manual selection process ensures a balanced distribution of images across subject classes and between photorealistic and non-photorealistic styles. In Fig. 5(a), we visualize the t-SNE [101, 77] of images from DreamBench and DreamBench++, which demonstrates the superiority of DreamBench++ in diversity. Besides, Fig. 5(b) presents the detailed statistics of the image distribution in DreamBench++.
3 Experiments
In this section, we provide evidence that DreamBench++ can more accurately reflect the capabilities of existing personalized image generation methods with its diverse and comprehensive data set. Additionally, we demonstrate that our systematic prompts-based GPT is a superior evaluator compared to existing metrics by analyzing their degree of alignment with human ratings. In order to further facilitate benchmarking research with GPT models, we conduct a comprehensive ablation study to show the designing key points for more human-aligned automated GPT evaluation.
Table 1: Evaluation of personalized image generation models on DreamBench++. All scores are normalized to 0-1, and -I & -T represent image & text, respectively.
Method T2I Model Concept Preservation Prompt Following
Human GPT DINO-I CLIP-I Human GPT CLIP-T
[…]
3.1 Experimental Setup
[…]
Figure 6: Comparison between images of high DINO score and high GPT-4o score. Instances with high human scores are ticked, and those with low human scores are crossed. DINO tends to yield high scores to images that preserve overall shape but do not put much weight on color, texture, and facial features, leading to frequent contradiction with human preference.
Human Annotators  We employ 7 human annotators to score for each instance in DreamBench++ to obtain ground truth human preference data. We provide human annotators with sufficient training to ensure they fully understand the personalized image generation task and can provide unbiased and discriminating scores. The scoring task and scheme given to humans are identical to those used for GPT, as described in Section 2. The GPT results and human results are isolated to avoid hindsight bias. Additionally, we ensure that each instance is rated by at least two humans to reduce noise.
Figure 7: A qualitative study of different methods on DreamBench++. We demonstrate the generation quality of different methods on our DreamBench++, including animals, humans, objects, and style, with photo and non-photo-realistic examples. The blue block highlights fine-tuning-based methods, and the green block highlights encoder-based methods. Instances above the dotted line are evaluated for subject preserving, and instances below are evaluated for style preserving.
3.2 Main Results
Quantitative Analysis  Table 1 shows the overall evaluation results, including human and GPT-4o rating scores. From the results, we observe that: i) DreamBench++ aligns better with humans than DINO or CLIP models. Driven by our dedicatedly-designed prompts, GPT-4o used by DreamBench++ yields impressive alignment with humans. This is because humans and DreamBench++ are all advanced in evaluating facial and textural characters and producing scores with a balanced consideration of all aspects. ii) DINO-I and CLIP-I yield significant divergence from humans in evaluating concept preservation. This could be because DINO/CLIP scores show a preference for images that preserve shapes or overall styles, as shown in Fig. 6. For example, methods proficient in posture and detailed texture transformation, such as DreamBooth-L, rank lower than humans and DreamBench++. iii) Traditional CLIP-T scores are as effective as DreamBench++ in evaluating prompt following, showing strong alignment with humans.
[…]
Animal Human Object Style Overall Photorealistic Style Imaginative Overall
[…]
Qualitative Analysis  Fig. 7 shows selected qualitative results on DreamBench++, which helps provide an intuitive understanding of evaluated models. With a more comprehensive and diverse collection of images, we have discovered numerous intriguing characteristics of these generation methods that were not apparent on existing datasets such as DreamBench. Specifically, we observe that: i) Fine-tuning-based methods outperform encoder-based methods on images containing more subject-oriented information, such as an animal or object, as they preserve more intricate details in the generated images. However, for images containing a person, fine-tuning-based methods often fail to preserve facial and clothing features. This suggests that the personalized generation of human images is more demanding for visual concept preservation than for textual following capabilities, which is an advantage for encoder-based methods. ii) However, for style-oriented cases when subject details are less critical, encoder-based methods perform better than fine-tuning-based methods. This further highlights the strengths of encoder-based methods in that they are more adept at recognizing and extracting high-level visual semantics, including overall shape, style, and thematic features.
Leaderboard  Table 2 shows the detailed results of the leaderboard with respect to the concept and prompt categories defined in Section 2. Note that: i) the human category shows the lowest average score of 0.482, which is -0.204 lower than the highest average score of animal. This category is very challenging in terms of concept preservation because humans are very sensitive to facial details, and many works are conducted specifically on it [105, 111, 100, 114]. ii) The object is also a relatively difficult category because the characteristics of each same or different object are often very diverse. In contrast, animals within the same category often share a large degree of visual similarity. iii) There exists a negative correlation between concept preservation and prompt following. The primary aim of personalized T2I evolution is to identify the Pareto optimum that balances both factors.
Table 3: Ablation study of prompt design. H, G, D, and C represent Human, GPT-4o, DINO Score, and CLIP Score, respectively. H-H value is also calculated to illustrate human self-alignment.
Method TI [28] DreamBooth [86] DreamBooth-L [86] BLIP-D [55] Emu2 [95] IP-Adapt.-P [115] IP-Adapt. [115]
T2I Model SD v1.5 SD v1.5 SDXL v1.0 SD v1.5 SDXL v1.0 SDXL v1.0 SDXL v1.0
Concept Preservation
H-H 0.685 0.647 0.656 0.613 0.746 0.602 0.591
G-H 0.544$\pm$0.014 0.596$\pm$0.003 0.641$\pm$0.007 0.362$\pm$0.017 0.669$\pm$0.005 0.366$\pm$0.017 0.458$\pm$0.002
- Internal Thinking -0.040 -0.023 -0.012 +0.001 -0.045 -0.038 -0.008
- Scoring Criteria -0.125 -0.116 -0.093 -0.158 -0.103 -0.227 -0.166
- Scoring Range -0.038 -0.017 -0.027 -0.006 -0.016 -0.009 -0.017
+ Human Prior -0.033 -0.022 -0.006 -0.015 -0.022 +0.009 -0.019
+ GPT4V -0.105 -0.067 -0.131 -0.180 -0.016 -0.301 -0.250
Prompt Following
H-H 0.475 0.516 0.469 0.619 0.441 0.576 0.509
G-H 0.461$\pm$0.007 0.506$\pm$0.002 0.402$\pm$0.001 0.541$\pm$0.003 0.422$\pm$0.011 0.484$\pm$0.006 0.531$\pm$0.006
- Internal Thinking -0.013 +0.004 -0.032 -0.002 -0.014 +0.012 -0.002
- Scoring Criteria -0.025 -0.012 -0.009 -0.012 -0.018 -0.017 -0.013
- Scoring Range -0.010 -0.013 -0.011 +0.043 -0.038 +0.060 +0.036
+ GPT4V -0.010 +0.012 $0$0.111 -0.007 -0.161 -0.134
3.3 Ablation Study
Table 3 shows the ablation study of the prompt design influences on alignment with humans. From the results, we observe that: i) The proposed prompt designs are all necessarily effective, demonstrating the superiority of the prompting method in DreamBench++. For example, removing the proposed internal thinking leads to a significant drop, indicating the effectiveness of self-alignment for alignment with humans. ii) The capability of the multimodal GPT used is scalable. This shows that DreamBench++ has the potential to be improved with stronger GPT models in the future. iii) Some human prior knowledge, such as reminding the GPT not to consider background when assessing visual concept preservation, leads to performance degradation.
Table 4: The human alignment degree among different evaluation metrics, measured by Krippendorff’s alpha value. H, G, D, and C represent Human, GPT-4o, DINO Score, and CLIP Score, respectively. H-H value is also calculated to illustrate human self-alignment.
Method T2I Model Concept Preservation $\bm{\mathrm{Kd}_{\bar{O}}}$ Prompt Following $\bm{\mathrm{Kd}_{\bar{O}}}$
H-H G-H D-H C-H H-H G-H C-H
[…]
$\bm{\mathrm{Ratio}_{\bar{O}}}$ 100% 79.64% 25.54% 3.34% 100% 93.18% 42.48%
4 Discussions
4.1 Is DreamBench++ aligned with humans?
Table 4 shows a more rigorous study of human alignment level using the mean Krippendorff’s alpha value [34]. The results show that DreamBench++ is a highly human-aligned benchmark. Notably, DreamBench++ achieves 79.64% and 93.18% evaluation consistency with human’s evaluation in concept preservation and prompt following capabilities, respectively. This result is +54.1% and +50.7% higher than traditional DINO and CLIP metrics.
Table 5: DreamBench [86] and DreamBench++ results comparison with traditional metrics. Unlike DreamBench, DreamBench++ uses one single reference image per instance; thus, the training steps and learning rate of $\bullet$ fine-tuning-based methods are slightly reduced to avoid overfitting.
Method T2I Model DreamBench [86] DreamBench++
DINO-I CLIP-I CLIP-T DINO-I CLIP-I CLIP-T
[…]
4.2 Is data diversity necessary?
To investigate the necessity of diverse and rich evaluation data, we compare the results on DreamBench and DreamBench++ using DINO and CLIP metrics. Table 5 demonstrates that diverse data in DreamBench++ is necessary to avoid biased evaluation. Although the results are overall consistent, especially for encoder-based methods, fine-tuning-based methods such as TI and DreamBooth both show a notable score drop. We argue that this is due to the reduced training steps and learning rate limited by the single-image fine-tuning, which may make the model not fully converged. However, unlike other encoder-based methods, Emu2 shows a significant score drop. Emu2 performs well only on natural images and simple text, while it cannot generate valid results when texts are complex or stylized or with anime reference images, as shown in Fig. 8.
Figure 8: Case study of successful and failure case on DreamBench++ with Emu2.
Table 6: Study of Chain-of-Though (CoT) and In-context Learning (ICL) on human alignment.
Method TI [28] DreamBooth [86] DreamBooth-L [86] BLIP-D [55] Emu2 [95] IP-Adapt.-P [115] IP-Adapt. [115]
T2I Model SD v1.5 SD v1.5 SDXL v1.0 SD v1.5 SDXL v1.0 SDXL v1.0 SDXL v1.0
H-H 0.685 0.647 0.656 0.613 0.746 0.602 0.591
w/o CoT & w/o ICL 0.544 0.596 0.641 0.362 0.669 0.366 0.458
+ 1 shot ICL -0.046 -0.019 -0.043 +0.013 -0.028 -0.098 -0.066
+ 2 shot ICL -0.042 -0.023 -0.022 -0.033 -0.036 -0.085 -0.054
w/ CoT & w/o ICL 0.510 0.576 0.602 0.329 0.644 0.359 0.418
+ 1 shot ICL -0.040 -0.008 -0.009 -0.020 -0.035 -0.145 -0.086
+ 2 shot ICL -0.030 -0.002 -0.002 -0.051 -0.031 -0.155 -0.082
4.3 Can we use free lunch to improve DreamBench++ evaluation?
Table 6 shows the result of utilizing free lunch techniques, including chain-of-thought (CoT) [108] and In-Context Learning (ICL) [4, 12]. CoT indicates that GPT-4 articulates its reasoning process before scoring, and ICL indicates GPT-4o is provided with human-written few-shot examples.
Chain-of-Thought: i) CoT is effective in evaluating prompt following capability. Through CoT, the model more accurately discerns the significance of phrases such as “morphs into a mythical dragon”, allowing it to assign a more appropriate evaluation score. ii) CoT does not bring improvement in concept preservation evaluation. We argue that CoT may shift attention to unnecessarily important background or texture information, as shown in Fig. 9.
Figure 9: Case study on CoT Prompting. We find that (a) CoT prompting can improve text following evaluation by recognizing important parts of the prompt. (b) However, it may also hinder visual concept preservation by drifting GPT’s attention away from the subject.
In-Context Learning: ICL counterintuitively leads to a drop in alignment. This could be attributed to the patching scheme, sample selection, or inherent bias within GPT-4o, making it non-trivial to prompt effectively. Thus, we provide our detailed prompt and hope to inspire future works.
5 Related Works
Personalized Image Generation aims to preserve concept consistency while accommodating the diverse contexts suggested by the instructions. In general, it can be traced back to early efforts on pixel-to-pixel (Pix2Pix) translation where the personalization orientation is free-form texts [11, 99, 75] or predefined translation across styles, seasons, species, or plants, etc [46, 127, 118, 106, 87]. Modern efforts go beyond Pix2Pix translation toward a free-form image generation conditioned on both reference images and prompts. Some works focus on fine-tuning techniques that turn a general T2I model into a specialist personalization model [28, 86, 53, 92, 74] using LoRA [42] or contrastive learning [117, 35], learning the subject or style information by reconstructive autoencoding [102, 23]. However, the necessity to fine-tune for new subjects limits their scalability. In contrast, encoder-based methods can generate subject-guided or style-guided images or edit images following prompts with one shot. Encoder- or adapter-based methods [125, 115, 109, 55, 48, 29, 16, 103, 105] train an encoder to encode the conditional image into embeddings, which are integrated into cross-attention mechanism in the diffusion process [40, 93, 68, 94]. Adapter-free methods [66, 64, 37, 36, 11] extract the information, such as attention maps [36, RBModulation24] from reference images, and fuse them into the image generation process. Furthermore, multimodal large language models (MLLMs) that are trained on extensive multimodal sequences can also serve as general foundation models [24, 95, 73, 30].
Benchmarking Image Generation involves a variety of metrics that focus on different aspects. Inception Score [89] and FID [39] judge image quality, while LPIPS [120], DreamSim [26], CLIP-I [81], and DINO Score [13] measure perceptual similarity. In text-guided generation, prompt-image alignment can be assessed by CLIP-T [81], CLIPScore [38], and BLIP Score [56, 57]. However, these metrics often fall short of reflecting human perception. To address this, human-aligned metrics [52, 112, 54] have been introduced, offering a more perceptive evaluation. Yet, they face limitations in scaling with the pace of new model developments. Thus, the necessity for automated and sustainable evaluation methods has emerged, with some [112, 60, 32] leveraging reward-model-based methods to encode human preferences, while others [51, 18, 110, 121, 43, 65] use multimodal  [12, 84, 5, 62] to automate the process and better mirror human tastes. While MLLM-based methods show promise in aligning with human preferences [121, 110, 45, 19], automated personalized evaluation remains an unresolved issue. VIEScore [51] assesses image generation quality by prompting GPT-4V [70] and LLaVA [62, 61], but is limited to four models in subject-driven tasks and obtains suboptimal results. Meanwhile, Dreambench [86], a common benchmark for personalized generative evaluation, only consists of 30 simple objects and lacks diversity comprehensiveness.
6 Conclusions
This paper introduces DreamBench++, a human-aligned personalized image generation benchmark. Extensive and comprehensive experiments have demonstrated our advantages in dataset diversity and complexity, as well as the alignment of our evaluation metrics with human preferences. In addition, we offer insights into prompt design for advanced multimodal GPTs, emphasizing the potential and challenges of enhancing GPT evaluation through chain-of-thought prompting and in-context learning. Our work aims to support future research on personalized image generation by providing a human-aligned benchmark and heuristics in utilizing advanced multimodal GPTs in visual evaluation.
Appendix A Implementation Details
Table 7: Training hyperparameters on DreamBench and DreamBench++. BS: batch size, LR: learning rate, Steps: training steps.
DreamBench DreamBench++
Method T2I Model BS LR Steps BS LR Steps
Textual Inversion [28] SD v1.5 4 5e-4 3000 1 5e-4 3000
Dreambooth [86] SD v1.5 1 2.5e-6 1000 1 2.5e-6 250
Dreambooth LoRA [86] SDXL v1.0 4 5e-5 500 1 5e-5 500
The configurations for the training hyperparameters used in training-based methods on DreamBench and DreamBench++, are detailed in Table 7. During the inference stage, all methods employ a guidance_scale of 7.5 and execute 100 inference steps, with the exception of Emu2, which uses a guidance_scale of 3 and performs 50 inference steps. Furthermore, BLIP-Diffusion and IP-Adapter incorporate negative prompts, as demonstrated in Table 8. Specifically, IP-Adapter includes an additional parameter, ip_adapter_scale, set at 0.6.
Table 8: Negative Prompt Templates
Method T2I Model Negative Prompt
BLIP-Diffusion [55] SD v1.5 over-exposure, under-exposure, saturated, duplicate, out of frame, lowres, cropped,
worst quality, low quality, jpeg artifacts, morbid, mutilated, ugly, bad anatomy,
bad proportions, deformed, blurry, duplicate
IP-Adapter ViT-G [115] SDXL v1.0 deformed, ugly, wrong proportion, low res, bad anatomy, worst quality, low quality
IP-Adapter-Plus ViT-H [115] SDXL v1.0 deformed, ugly, wrong proportion, low res, bad anatomy, worst quality, low quality
We dedicate significant effort to tuning hyperparameters to ensure that the performance of each method on DreamBench is consistent with results reported in original papers. As shown in Table 9, the results of our reproduction are comparable to or even better than the official results reported by the original papers.
Table 9: Reproduced results. Our reproduction is comparable to or better than the official results. N/A denotes that the official paper does not report the corresponding results.
Method T2I Model DINO-I CLIP-I CLIP-T
Official Reproduction Official Reproduction Official Reproduction
[…]
Appendix B Limitation & Future Work
Human-aligned evaluation & benchmarking is an emerging but challenging direction, and we have only made preliminary attempts at personalized image generation. Moreover, our evaluation results heavily rely on the advancements of multimodal large language models and require carefully designed system prompts. We believe that as visual world models continue to develop, the evaluation performance will be further optimized. Our future work will focus on more applications with human-aligned evaluation, such as 3D generation [78, 79, 63, 27], video generation [41, 91, 9], autonomous driving [33, 58, 119], and even embodied visual intelligence [59, 10, 25, 80, 123, 31].
Broader Impact
Powerful as the T2I generative models pretrained on large-scale web-scraped data, the models may be misused as illegal or unethical tools for generating NSFW content. This potential impact can also be brought by personalized T2I models as they are typically built on the pretrained T2I foundation models. As a result, it is critical to use tools such as NSFW detectors to avoid such content during both usage and evaluation. For example, the data used for evaluation must avoid the NSFW content by data filtering. In this paper, such contents are filtered out by human annotators.

### v2

@@
 Abstract
-Personalized image generation holds great promise in assisting humans in everyday work and life due to its impressive function in creatively generating personalized content.
+Personalized image generation holds great promise in assisting humans in everyday work and life due to its impressive ability to creatively generate personalized content across various contexts.
 However, current evaluations either are automated but misalign with humans or require human evaluations that are time-consuming and expensive.
-In this work, we present DreamBench++, a human-aligned benchmark automated by advanced multimodal GPT models.
+In this work, we present DreamBench++, a human-aligned benchmark that advanced multimodal GPT models automate.
 Specifically, we systematically design the prompts to let GPT be both human-aligned and self-aligned, empowered with task reinforcement.
@@
 By benchmarking 7 modern generative models, we demonstrate that DreamBench++ results in significantly more human-aligned evaluation, helping boost the community with innovative findings.
-Figure 1: Overview of DreamBench++. We collect diverse images and prompts, and utilize GPT-4o for automated evaluation aligned with human preference.
+Figure 1: Overview of DreamBench++.
+We collect diverse images and prompts as the base dataset and utilize GPT-4o for automated evaluation aligned with human preference.
+To evaluate personalized image generation, DreamBench++ mainly focuses on image & image concept preservation and text & image prompt following assessments.
 1 Introduction
-Driven by the significant advances in large-scale text-to-image (T2I) generative models [85, 82, 8, 83, 69, 88, 116, 14, 27, 21, 22, 7, 49, 24], it is now possible to generate images conditioned on not only arbitrary text prompts but also by given reference images—personalized image generation [86, 28, 55, 115, 53, 29, 6, 17, 48, 15, 111, 98, 109, 67, 44, 105, 66, 103, 16, 99, 126].
-In general, to be useful as an artistic creation tool for inspiration or products [113], the following two basic criteria must be fulfilled: i) Prompt following (image & prompt consistency).
-Generated images must follow the prompt description, which is a requirement shared with vanilla T2I generation [8, 83].
+Driven by the significant advances in large-scale text-to-image (T2I) generative models (Rombach et al., 2022; Ramesh et al., 2021; Betker et al., 2023; Ramesh et al., 2022; Nichol et al., 2022; Saharia et al., 2022b; Yu et al., 2022; Chang et al., 2023; Gafni et al., 2022; Ding et al., 2021; 2022; Balaji et al., 2022; Kang et al., 2023; Dong et al., 2024), it is now possible to generate images conditioned on not only arbitrary text prompts but also by given reference images—personalized image generation (Ruiz et al., 2023; Gal et al., 2023a; Li et al., 2023a; Ye et al., 2023; Kumari et al., 2023; Gal et al., 2023b; Arar et al., 2023; Chen et al., 2023c; Jia et al., 2023; Chen et al., 2023a; Xiao et al., 2023; Tewel et al., 2023; Wei et al., 2023; Ma et al., 2023; Hua et al., 2023; Wang et al., 2024b; Lv et al., 2024; Wang et al., 2024a; Chen et al., 2023b; Tumanyan et al., 2023; Zhou et al., 2024; Tan et al., 2024; He et al., 2024c; Wang et al., 2024c; Wu et al., 2024a; He et al., 2024a; Xiao et al., 2024; Arar et al., 2024; Huang et al., 2024b; a; Pang et al., 2024a; b; Qiu et al., 2023; Hu et al., 2024).
+In general, to be useful as an artistic creation tool for inspiration or products (Yacoubian, 2022; Xie et al., 2024), the following two basic criteria must be fulfilled: i) Prompt following (image & prompt consistency).
+Generated images must follow the prompt description, which is a requirement shared with vanilla T2I generation (Betker et al., 2023; Ramesh et al., 2022).
 ii) Concept preservation (image & image consistency).
-For personalized image generation, the concept of the reference image, i.e., the main subject’s semantic details (e.g., facial characters) or high-level abstractions (e.g., overall style), must be preserved in the generated image.
-For example, a user may want to “imagine his own dog traveling around the world” [86], and the generated dog must be the same as his but traveling.
-Figure 2: Qualitative comparison of concept preservation evaluation between DreamBench++ and traditional DINO [13].
+For personalized image generation, the concept of the reference image, i.e., the main subject’s semantic details (e.g., facial characters) or high-level abstractions (e.g., overall style), must be preserved.
+For example, a user may want to “imagine his own dog traveling around the world” (Ruiz et al., 2023), and the generated dog must be the same as his but traveling.
+Figure 2: Qualitative comparison of concept preservation evaluation between DreamBench++ and traditional DINO (Caron et al., 2021).
 DINO often fails to yield human-aligned evaluation while our DreamBench++ succeeds using multimodal GPT models as the evaluator.
 To meet the aforementioned requirements, numerous efforts have been devoted.
-One line of fine-tuning-based works focuses on fine-tuning general T2I models to specialist personalization models by reproducing specific concepts present in training sets [86, 28, 17, 53].
-Meanwhile, another line of encoder-based works, instead, achieves concept-preservation by training features adaptation to inject reference image features into a general T2I model [115, 29, 6, 24, 95, 96, 73].
-Although remarkable results have been achieved, one question arises: can we comprehensively evaluate these models to figure out which technical route is superior and where to head?
+One line of fine-tuning-based works focuses on fine-tuning general T2I models to specialist personalization models by reproducing specific concepts present in training sets (Ruiz et al., 2023; Gal et al., 2023a; Chen et al., 2023c; Kumari et al., 2023).
+Meanwhile, another line of encoder-based works, instead, achieves concept-preservation by training features adaptation to inject reference image features into a general T2I model (Ye et al., 2023; Gal et al., 2023b; Arar et al., 2023; Dong et al., 2024; Sun et al., 2024a; b; Pan et al., 2024).
+Despite remarkable progress, one question arises: can we comprehensively evaluate these models to figure out which technical route is superior and where to head?
+Figure 3: Overall procedure of prompting GPT-4o for automated evaluation.
+Human-written meta-prompts provide task description, scoring criteria, range, and format.
+GPT-4o then follows reasoning instructions for self-aligned task summarization and planning.
+Finally, all prompts and reasoning outputs are combined with image samples for score generation.
 In this work, we aim to answer this question by developing a new benchmark that properly evaluates personalized T2I models driven by the above two requirements.
@@
     Human-Aligned  As shown in Fig.
-2, traditional metrics like DINO [13] and CLIP [81] often result in significant discrepancies from humans.
-This is caused by the image similarity measurement nature of DINO and CLIP models, and thus crowd-sourced human evaluation is typically necessary for obtaining a correct quantitative understanding of generated images [54, 52, 112].
+2, traditional metrics like DINO (Caron et al., 2021) and CLIP (Radford et al., 2021) often result in significant discrepancies from humans.
+This is caused by the image similarity measurement nature of DINO and CLIP models, and thus crowd-sourced human evaluation is typically necessary for obtaining a correct quantitative understanding of generated images (Lee et al., 2023; Ku et al., 2024; Xu et al., 2023).
 Therefore, different from existing works that utilize CLIP and DINO as metrics that may be humanly misaligned, our DreamBench++ demonstrates surprisingly consistent evaluation results aligned with humans.
 For instance, by evaluating 7 modern models, DreamBench++ achieves 79.64% and 93.18% agreement with human’s evaluation in concept preservation and prompt following capabilities, respectively.
-Notably, it is +54.1% and +50.7% higher than traditional DINO and CLIP metrics.
+Notably, it is +32.59% and +37.23% higher than traditional DINO and CLIP metrics.
   • 2.
-    Automated  However, it is non-standardized and expensive to perform high-quality human evaluations, making human evaluations non-trivial to use.
-To address this challenge, DreamBench++ achieves automated but human-aligned evaluation by using advanced multimodal GPT models as metrics such as GPT-4o [71].
+    Automated  However, it is non-standardized and expensive to perform high-quality human evaluations.
+To address this challenge, DreamBench++ achieves automated but human-aligned evaluation by using advanced multimodal GPT models such as GPT-4o (OpenAI, 2024) as metrics.
 The challenges lie in two aspects: i) prompt design and ii) reasoning procedure for scoring.
 We systematically standardize the automated GPT evaluation by first designing the evaluation instructions that provide overall task requirements, where language is a general interface for instructing human preference.
-Inspired by Self-Align [97], we instruct the GPT model to conduct internal thinking that aligns itself for better task and preference understanding.
-Then, the GPT model provides the summary & planning for the task and scoring criteria, and the final scores are provided with optional chain-of-thought (CoT) [108, 122].
+Inspired by Self-Align (Sun et al., 2023), we instruct GPT to conduct internal thinking that aligns itself for better task and preference understanding.
+Then, GPT provides the summary & planning for the task and scoring criteria, and the final scores are provided with optional chain-of-thought (CoT) (Wei et al., 2022; Zhang et al., 2023d).
   • 3.
-    Diverse  To avoid biased understanding due to the low diversity of evaluation data, DreamBench++ collects a large number of diverse images covering various levels of difficulty categories, from animals and styles that are relatively simple to more challenging human subjects, objects, and non-natural image styles (See Fig.
+    Diverse  To avoid bias from low-diversity evaluation data, DreamBench++ compiles a wide range of images, covering varying levels of difficulty from simpler animals and styles to more complex human subjects, objects, and non-natural styles (see Fig.
 1).
-Compared to commonly used DreamBench [86] that only contains 30 subjects and 25 prompts, DreamBench++ scales up the benchmarking data to 150 images and 1,350 prompts, i.e., 5$\times$ and 54$\times$ more than DreamBench.
-As a result, when evaluating models on this larger dataset, different but more robust conclusions have been obtained that demonstrate an unbiased and more comprehensive evaluation of DreamBench++.
-Takeaways  Here, we offer some insightful findings that we obtain by evaluating 7 modern personalized T2I models: i) DINO-based rating puts tremendous weight on resemblance in overall shape or color and neglects detailed visual features, which makes it a sub-optimal metric for personalized image generation evaluation; ii) The main goal of personalized image generation is to find the Pareto optimal solution that balances concept preservation and prompt following.
-Among these 7 models, DreamBooth [86] is the best in overall performance, preserving highly detailed visual features while keeping close adherence with text prompt; iii) Currently, personalized T2I models perform well on animal and style category.
-However, they have not performed well in the human category due to sensitivity to facial details and in the object category due to its diversity.
-Existing works [105, 100, 114, 115, 111] focuses on preserving facial features to address the former problem, but the latter has not been extensively explored.
+Unlike DreamBench (Ruiz et al., 2023), which includes only 30 subjects and 25 prompts, DreamBench++ significantly expands the dataset to 150 images and 1,350 prompts—5$\times$ and 54$\times$ more, respectively.
+While CustomConcept101 (Kumari et al., 2023) offers 101 subjects, its diversity is limited by repetitive image categories and a focus on photorealistic styles, with simple prompts that restrict its ability to evaluate models on more complex tasks.
+Consequently, DreamBench++ enables more robust and comprehensive conclusions in model evaluation.
+Takeaways  We present some insightful findings from evaluating seven modern personalized T2I models: i) DINO-based ratings prioritize overall shape and color over detailed features, making them suboptimal for evaluating personalized image generation; ii) The primary goal is to achieve a Pareto optimal balance between concept preservation and prompt adherence.
+Among the models, DreamBooth (Ruiz et al., 2023) excels in preserving detailed visual features while closely following text prompts; iii) Current models perform well in animal and style categories but struggle with human images due to sensitivity to facial details and diverse object categories.
+While existing work (Wang et al., 2024b; Valevski et al., 2023; Yan et al., 2023; Ye et al., 2023; Xiao et al., 2023) addresses facial feature preservation, the challenge of object diversity remains underexplored.
 We are presenting DreamBench++ with open-sourced codes and evaluation standardization to promote innovation within the research community.
@@
 2 DreamBench++
-We introduce DreamBench++, a human-aligned, automated, and diverse benchmark that evaluates the two capabilities of personalized image generation models.
-In the following, we describe how we construct DreamBench++ from two aspects: prompts and data.
-Figure 3: Overall procedure of prompting GPT-4o for automated evaluation.
-The evaluation instructions are meta-prompting information written by humans, including task description, scoring criteria, scoring range, and format specification.
-Then, GPT-4o is prompted with reasoning instructions to perform internal thinking that provides a self-aligned task summary and planning.
-Finally, all prompts and reasoning outputs are joined with image samples for score outputs.
-2.1 Prompting Multimodal GPT for Automated & Human-Aligned Benchmarking
-It is challenging to obtain a solid quantitative understanding of generated models, especially when evaluating visual contents like images that rely on human evaluations [54, 52].
-Thus, it is critical to achieve automated evaluation by utilizing multimodal GPT models, which are trained particularly in the principle of aligning with human preference [72, 20, 71, 70].
-This is evidenced by the recent progress achieved by Wu et al., which demonstrates that GPT-4V [70] can serve as a human-aligned text-to-3D generation evaluator.
+2.1 Prompting GPT for Automated & Human-Aligned Benchmarking
+It is challenging to obtain a solid quantitative understanding of generated models, especially when evaluating visual contents that rely on human evaluations (Lee et al., 2023; Ku et al., 2024).
+Thus, it is critical to achieve automated evaluation by utilizing multimodal GPT models, which are trained particularly in the principle of aligning with human preference (Ouyang et al., 2022; Christiano et al., 2017; OpenAI, 2024; 2023).
+This is evidenced by the recent progress achieved by Wu et al., which demonstrates that GPT-4V (OpenAI, 2023) can serve as a human-aligned text-to-3D generation evaluator.
 However, as pointed out by Zhang et al.
 and Ku et al., multimodal GPT models often fall short in evaluating personalized image generation—often more challenging when distinguishing subtle difference for concept preservation assessment using GPT—still underexplored.
-To tackle this issue, we detail how we systematically design the prompt of multimodal GPT (GPT-4o [71], by default) for human alignment reinforcement but also improve the reasoning progress that helps the GPT models to be more self-aligned, introduced as follows.
+To tackle this issue, we detail how we systematically design the prompt of multimodal GPT (GPT-4o (OpenAI, 2024), by default) for human alignment reinforcement but also improve the reasoning progress that helps the GPT models to be more self-aligned, introduced as follows.
 Compare or rate?
-There are typically two schemes for quantitatively evaluating generative models in human evaluations: rating and comparison [121, 124].
+There are typically two schemes for quantitatively evaluating generative models in human evaluations: rating and comparison (Zhang et al., 2023c; Zheng et al., 2023).
 The rating scheme requires human reviewers to assign an absolute score to each instance, while the comparison scheme asks human reviewers to express a relative preference among different instances.
 Though effective as the comparison scheme is when humans are involved, we find that there are two critical issues.
-i) Positional Bias: the scoring results of GPT-4V/GPT-4o is sensitive to the order in which images are presented [70, 104, 107, 121, 110, 124], making it unsuitable for comparison scheme.
+i) Positional Bias: the scoring results of GPT-4V/GPT-4o is sensitive to the order in which images are presented (OpenAI, 2023; Wang et al., 2023a; b; Zhang et al., 2023c; Wu et al., 2024b; Zheng et al., 2023), making it unsuitable for comparison scheme.
 ii) Quadratic Complexity: As the number of methods increases, the number of essential evaluation runs for numerical rating increases linearly, while the number of comparative assessments increases quadratically.
@@
 Figure 4: Dataset construction process of our DreamBench++.
-We start by obtaining keywords through GPT generation, existing datasets, and human proposals.
-Next, we collect corresponding images from the internet.
-These images are then filtered to remove low-quality ones through both model and human assessment.
-The remaining high-quality images are used as input for GPT-4o to generate text prompts of varying difficulty levels.
+We generate keywords using GPT-4, existing datasets, and human input, then collect corresponding web images.
+After filtering out low-quality images via model and human assessment, the high-quality ones are used as input for GPT-4o to generate text prompts of varying difficulty.
 Evaluation Instructions  The evaluation instructions serve as the meta-prompting that describes overall tasks, which is shown in Fig.
@@
 Reasoning Instructions  Given the evaluation instructions, it is crucial to reinforce the alignment with both the human instruction and itself to largely leverage the pretrained knowledge.
-To this end, we adopt a 2-step evaluation policy as follows: i) Internal Thinking: Inspired by Self-Align [97], we introduce internal thinking to strengthen task understanding and instruction following capabilities.
+To this end, we adopt a 2-step evaluation policy: i) Internal Thinking: Inspired by Self-Align (Sun et al., 2023), we introduce internal thinking to strengthen task understanding and instruction following capabilities.
 Specifically, we prompt the GPT model by asking if it understands the task or not and let it summarize the task.
 ii) Summary & Planning: According to the given internal thinking instruction, the GPT will summarize and plan for the evaluation task itself.
-It can also be viewed as a generalized form of chain-of-thought reasoning [108, 122].
-The complete procedure is illustrated in Fig.
+It can also be viewed as a generalized form of chain-of-thought reasoning (Wei et al., 2022; Zhang et al., 2023d).
+See Fig.
 3.
 2.2 Scaling Up Personalized Image Generation Benchmarking
-Pioneering works like DreamBooth [86] and SuTI [17] have successfully set up baseline datasets for the evaluation of personalized image generation, and DreamBench++ follows them to categorize images into three types: ❶ objects, ❷ living subjects, and ❸ styles.
-However, due to the small-scale nature of DreamBench, it is limited as some methods may converge well on its samples while performing unsatisfactorily on other data.
-To avoid this possible biased evaluation, we scale up the benchmarking data by increasing both image numbers and diversity.
-Figure 5: Data distribution visualization. (a) Images comparison between DreamBench and DreamBench++ using t-SNE [101, 77]. (b) Image and prompt distribution of DreamBench++.
-Data Construction from Internet  There are broad images on the Internet, and many datasets are constructed from it [90, 47].
-DreamBench++ mainly collects images from Unsplash [3], Rawpixel [2], and Google Image Search [1], along with contributions from individuals with authorized permissions.
+Pioneering works like DreamBooth (Ruiz et al., 2023), SuTI (Chen et al., 2023c) and CustomConcept101 (Kumari et al., 2023) have successfully set up baseline datasets evaluating personalized image generation, and DreamBench++ follows them to categorize images into three types: ❶ objects, ❷ living subjects, and ❸ styles.
+However, DreamBench’s small scale and CustomConcept101’s limited diversity may lead to biased evaluations, as some methods may converge well on its samples while performing unsatisfactorily on other data.
+To mitigate this, we scale up data by increasing both the number and diversity of images.
+Figure 5: Data distribution visualization. (a) Images comparison between DreamBench and DreamBench++ using t-SNE. (b) Image and prompt distribution of DreamBench++.
+Data Construction from Internet  The internet offers vast images, with many datasets built from it (Schuhmann et al., 2021; Jia et al., 2021).
+DreamBench++ primarily sources images from Unsplash (uns, ), Rawpixel (raw, ), and Google Image Search (goo, ), along with authorized individual contributions.
 Each image’s copyright status has been verified for academic suitability.
 As shown in Fig.
-4, we collect and construct high-quality data in DreamBench++ by following 3 steps:
+4, we collect and curate high-quality data to ensure robustness and diversity.
   • Keywords Generation  First, we generate 200 relevant keywords using GPT-4o and join them with the 200 most frequent keywords from Unsplash.
 After filtering out duplicated keywords, seven human annotators will extend the list to around 300 based on their interests.
-  • Internet Images Collection  Given selected keywords, we retrieved corresponding images from Unsplash, Rawpixel, and Google Image Search.
-To filter out images unsuitable for personalized image generation, SAM [50] is applied to identify subject regions in images and discard those with too small subject areas.
-Human annotators will then filter out images with noisy backgrounds.
-Curated images were cropped to centralize the subject, resulting in two images per keyword.
-Keywords that fail to yield suitable images will be discarded in this process.
+  • Internet Images Collection  Given selected keywords, we retrieved images from Unsplash, Rawpixel, and Google Image Search.
+To filter out unsuitable images, SAM (Kirillov et al., 2023) identifies subject regions, discarding those with small subject areas.
+Human annotators remove images with noisy backgrounds.
+Curated images are then cropped to centralize the subject, yielding two images per keyword.
+Keywords without suitable images are discarded.
   • Prompt Generation  After image collection, 9 text prompts per image were generated using GPT-4o, designed to cover a range of difficulties: 4 prompts for ❶ photorealistic styles, 3 for ❷ non-photorealistic styles, and 2 for ❸ complicated & imaginative contents.
-To align with established evaluation methods, we use few-shot prompts selected from PartiPrompts [116].
+To align with established evaluation methods, we use few-shot prompts selected from PartiPrompts (Yu et al., 2022).
 Human calibration ensures that all generated prompts are ethical and without flaws.
-As a result, the construction process finally yields 150 high-quality images and 1,350 prompts.
-Diversity Visualization  When collecting images, we notice that Internet images possess a bias towards photorealistic styles.
+As a result, the construction process finally yielded 150 high-quality images and 1,350 prompts.
+Diversity Visualization  Internet images are numerous.
+However, there is a bias towards photorealistic styles.
 To diversify, various non-photorealistic styles are enlisted, and human annotators are tasked to gather images for each style, including anime, sketches, traditional Chinese paintings, artworks, and cartoon characters from games.
-Then, a manual selection process ensures a balanced distribution of images across subject classes and between photorealistic and non-photorealistic styles.
+Then, a manual selection process ensures a balanced distribution across subject classes and between photorealistic and non-photorealistic styles.
 In Fig.
-5(a), we visualize the t-SNE [101, 77] of images from DreamBench and DreamBench++, which demonstrates the superiority of DreamBench++ in diversity.
+5(a), we visualize the t-SNE (Van der Maaten & Hinton, 2008; Poličar et al., 2019) of images from DreamBench and DreamBench++, which demonstrates the superiority of DreamBench++ in diversity.
 Besides, Fig.
-5(b) presents the detailed statistics of the image distribution in DreamBench++.
+5(b) presents the detailed image distribution in DreamBench++.
 3 Experiments
-In this section, we provide evidence that DreamBench++ can more accurately reflect the capabilities of existing personalized image generation methods with its diverse and comprehensive data set.
-Additionally, we demonstrate that our systematic prompts-based GPT is a superior evaluator compared to existing metrics by analyzing their degree of alignment with human ratings.
-In order to further facilitate benchmarking research with GPT models, we conduct a comprehensive ablation study to show the designing key points for more human-aligned automated GPT evaluation.
 Table 1: Evaluation of personalized image generation models on DreamBench++. All scores are normalized to 0-1, and -I & -T represent image & text, respectively.
@@
 DINO tends to yield high scores to images that preserve overall shape but do not put much weight on color, texture, and facial features, leading to frequent contradiction with human preference.
-Human Annotators  We employ 7 human annotators to score for each instance in DreamBench++ to obtain ground truth human preference data.
-We provide human annotators with sufficient training to ensure they fully understand the personalized image generation task and can provide unbiased and discriminating scores.
+Human Annotators  We employ 7 human annotators to score each instance in DreamBench++ to obtain ground truth human preference data.
+We provide human annotators with sufficient training to ensure they fully understand the personalized T2I generation task and can provide unbiased and discriminating scores.
 The scoring task and scheme given to humans are identical to those used for GPT, as described in Section 2.
@@
 Additionally, we ensure that each instance is rated by at least two humans to reduce noise.
-Figure 7: A qualitative study of different methods on DreamBench++.
-We demonstrate the generation quality of different methods on our DreamBench++, including animals, humans, objects, and style, with photo and non-photo-realistic examples.
-The blue block highlights fine-tuning-based methods, and the green block highlights encoder-based methods.
-Instances above the dotted line are evaluated for subject preserving, and instances below are evaluated for style preserving.
 3.2 Main Results
-Quantitative Analysis  Table 1 shows the overall evaluation results, including human and GPT-4o rating scores.
-From the results, we observe that: i) DreamBench++ aligns better with humans than DINO or CLIP models.
+Quantitative & Qualitative Analysis  Table 1 shows the overall evaluation results, including human and GPT-4o rating scores.
+The results show that: i) DreamBench++ aligns better with humans than DINO or CLIP models.
 Driven by our dedicatedly-designed prompts, GPT-4o used by DreamBench++ yields impressive alignment with humans.
-This is because humans and DreamBench++ are all advanced in evaluating facial and textural characters and producing scores with a balanced consideration of all aspects.
+This is because humans and DreamBench++ are all advanced in evaluating facial and textural characters and producing scores with a balanced consideration.
 ii) DINO-I and CLIP-I yield significant divergence from humans in evaluating concept preservation.
-This could be because DINO/CLIP scores show a preference for images that preserve shapes or overall styles, as shown in Fig.
-6.
-For example, methods proficient in posture and detailed texture transformation, such as DreamBooth-L, rank lower than humans and DreamBench++.
+This could be because DINO/CLIP scores show a preference for images that preserve shapes or overall styles (see Fig.
+6).
 iii) Traditional CLIP-T scores are as effective as DreamBench++ in evaluating prompt following, showing strong alignment with humans.
+See qualitative results in Appendix A for an intuitive understanding of evaluated models.
 […]
 […]
-Animal Human Object Style Overall Photorealistic Style Imaginative Overall
+Animal Human Object Style Overall Realistic Style Imaginative Overall
 […]
@@
 […]
-Qualitative Analysis  Fig.
-7 shows selected qualitative results on DreamBench++, which helps provide an intuitive understanding of evaluated models.
-With a more comprehensive and diverse collection of images, we have discovered numerous intriguing characteristics of these generation methods that were not apparent on existing datasets such as DreamBench.
-Specifically, we observe that: i) Fine-tuning-based methods outperform encoder-based methods on images containing more subject-oriented information, such as an animal or object, as they preserve more intricate details in the generated images.
-However, for images containing a person, fine-tuning-based methods often fail to preserve facial and clothing features.
-This suggests that the personalized generation of human images is more demanding for visual concept preservation than for textual following capabilities, which is an advantage for encoder-based methods.
-ii) However, for style-oriented cases when subject details are less critical, encoder-based methods perform better than fine-tuning-based methods.
-This further highlights the strengths of encoder-based methods in that they are more adept at recognizing and extracting high-level visual semantics, including overall shape, style, and thematic features.
-Leaderboard  Table 2 shows the detailed results of the leaderboard with respect to the concept and prompt categories defined in Section 2.
+Leaderboard  Table 2 shows the leaderboard results with respect to the concept and prompt categories defined in Section 2.
 Note that: i) the human category shows the lowest average score of 0.482, which is -0.204 lower than the highest average score of animal.
-This category is very challenging in terms of concept preservation because humans are very sensitive to facial details, and many works are conducted specifically on it [105, 111, 100, 114].
-ii) The object is also a relatively difficult category because the characteristics of each same or different object are often very diverse.
-In contrast, animals within the same category often share a large degree of visual similarity.
+This category is very challenging in terms of concept preservation because due to facial details, and many works are conducted specifically on it (Wang et al., 2024b; Xiao et al., 2023; Valevski et al., 2023; Yan et al., 2023).
+ii) The object is also a relatively difficult category due to object diversity.
+In contrast, animals within the same category often share a strong visual similarity.
 iii) There exists a negative correlation between concept preservation and prompt following.
-The primary aim of personalized T2I evolution is to identify the Pareto optimum that balances both factors.
+The primary aim of personalized T2I evolution is to identify the Pareto optimum that balances both.
+iv) The CP/PF ratio reflects the over-fitting issue.
+A higher ratio indicates excessive adherence to the reference at the cost of prompt alignment, while a lower ratio suggests better prompt adherence but weaker concept preservation.
 Table 3: Ablation study of prompt design. H, G, D, and C represent Human, GPT-4o, DINO Score, and CLIP Score, respectively. H-H value is also calculated to illustrate human self-alignment.
-Method TI [28] DreamBooth [86] DreamBooth-L [86] BLIP-D [55] Emu2 [95] IP-Adapt.-P [115] IP-Adapt. [115]
+Method TI DreamBooth DreamBooth-L BLIP-D Emu2 IP-Adapt.-P IP-Adapt.
 T2I Model SD v1.5 SD v1.5 SDXL v1.0 SD v1.5 SDXL v1.0 SDXL v1.0 SDXL v1.0
-Concept Preservation
+Concept Preservation $\bm{\mathrm{Kd}_{\bar{O}}}$
 H-H 0.685 0.647 0.656 0.613 0.746 0.602 0.591
@@
 + GPT4V -0.105 -0.067 -0.131 -0.180 -0.016 -0.301 -0.250
-Prompt Following
+Prompt Following $\bm{\mathrm{Kd}_{\bar{O}}}$
 H-H 0.475 0.516 0.469 0.619 0.441 0.576 0.509
@@
 + GPT4V -0.010 +0.012 $0$0.111 -0.007 -0.161 -0.134
-3.3 Ablation Study
-Table 3 shows the ablation study of the prompt design influences on alignment with humans.
-From the results, we observe that: i) The proposed prompt designs are all necessarily effective, demonstrating the superiority of the prompting method in DreamBench++.
-For example, removing the proposed internal thinking leads to a significant drop, indicating the effectiveness of self-alignment for alignment with humans.
+Abaltion Study  Table 3 shows the ablation study of the prompt design influences on alignment measured by mean Krippendorff’s alpha value (Hayes & Krippendorff, 2007).
+We observe that: i) The proposed prompt designs are all necessarily effective, demonstrating the superiority of the prompting method in DreamBench++.
+For example, removing the proposed internal thinking leads to a significant drop, indicating the effectiveness of self-alignment.
 ii) The capability of the multimodal GPT used is scalable.
-This shows that DreamBench++ has the potential to be improved with stronger GPT models in the future.
+This shows that DreamBench++ has the potential to be improved in the future.
 iii) Some human prior knowledge, such as reminding the GPT not to consider background when assessing visual concept preservation, leads to performance degradation.
-Table 4: The human alignment degree among different evaluation metrics, measured by Krippendorff’s alpha value.
+4 Discussions
+4.1 Is DreamBench++ aligned with humans?
+Table 4 shows a more rigorous study of human alignment level using the mean Pearson correlation value (Pearson, 1895).
+The results show that DreamBench++ is a highly human-aligned benchmark.
+Notably, DreamBench++ achieves 83.31% and 98.71% evaluation consistency with human’s evaluation in concept preservation and prompt following capabilities, respectively.
+This result is +32.59% and +37.23% higher than traditional DINO and CLIP metrics.
+Table 4: The human alignment degree among different evaluation metrics, measured by Pearson correlation value.
 H, G, D, and C represent Human, GPT-4o, DINO Score, and CLIP Score, respectively.
 H-H value is also calculated to illustrate human self-alignment.
-Method T2I Model Concept Preservation $\bm{\mathrm{Kd}_{\bar{O}}}$ Prompt Following $\bm{\mathrm{Kd}_{\bar{O}}}$
+Method T2I Model Concept Preservation $\bm{\mathrm{PC}_{\bar{O}}}$ Prompt Following $\bm{\mathrm{PC}_{\bar{O}}}$
 H-H G-H D-H C-H H-H G-H C-H
@@
 […]
-$\bm{\mathrm{Ratio}_{\bar{O}}}$ 100% 79.64% 25.54% 3.34% 100% 93.18% 42.48%
-4 Discussions
-4.1 Is DreamBench++ aligned with humans?
-Table 4 shows a more rigorous study of human alignment level using the mean Krippendorff’s alpha value [34].
-The results show that DreamBench++ is a highly human-aligned benchmark.
-Notably, DreamBench++ achieves 79.64% and 93.18% evaluation consistency with human’s evaluation in concept preservation and prompt following capabilities, respectively.
-This result is +54.1% and +50.7% higher than traditional DINO and CLIP metrics.
-Table 5: DreamBench [86] and DreamBench++ results comparison with traditional metrics.
-Unlike DreamBench, DreamBench++ uses one single reference image per instance; thus, the training steps and learning rate of $\bullet$ fine-tuning-based methods are slightly reduced to avoid overfitting.
-Method T2I Model DreamBench [86] DreamBench++
+$\bm{\mathrm{Ratio}_{\bar{O}}}$ 100% 83.31% 50.72% 60.98% 100% 98.17% 61.48%
+4.2 Is data diversity necessary?
+To assess the importance of diverse data, we compare results on DreamBench and DreamBench++ using DINO and CLIP metrics.
+Table 5 shows that the diverse data in DreamBench++ is key to unbiased evaluation.
+While overall results are consistent, TI, DreamBooth, and Emu2 show notable score drops.
+These methods perform well on natural images and simple text but struggle with complex or stylized prompts and anime references, see Fig.
+7.
+Table 5: DreamBench and DreamBench++ results comparison with traditional metrics.
+Unlike DreamBench, DreamBench++ uses a single reference image per instance; thus, the training steps and learning rate of $\bullet$ fine-tuning-based methods are slightly reduced to avoid overfitting.
+Method T2I Model DreamBench DreamBench++
 DINO-I CLIP-I CLIP-T DINO-I CLIP-I CLIP-T
@@
 […]
-4.2 Is data diversity necessary?
-To investigate the necessity of diverse and rich evaluation data, we compare the results on DreamBench and DreamBench++ using DINO and CLIP metrics.
-Table 5 demonstrates that diverse data in DreamBench++ is necessary to avoid biased evaluation.
-Although the results are overall consistent, especially for encoder-based methods, fine-tuning-based methods such as TI and DreamBooth both show a notable score drop.
-We argue that this is due to the reduced training steps and learning rate limited by the single-image fine-tuning, which may make the model not fully converged.
-However, unlike other encoder-based methods, Emu2 shows a significant score drop.
-Emu2 performs well only on natural images and simple text, while it cannot generate valid results when texts are complex or stylized or with anime reference images, as shown in Fig.
+Figure 7: Case study of successful and failure case on DreamBench++. The left images are reference images and the right images are results generated by Emu2, DreamBooth, and TI.
+4.3 Can we use free lunch to improve DreamBench++ evaluation?
+Table 6 shows the result of utilizing free lunch techniques, including chain-of-thought (CoT) (Wei et al., 2022) with GPT-4 articulating reasoning process and In-Context Learning (ICL) (Alayrac et al., 2022; Brown et al., 2020) with human-written few-shot examples.
+Chain-of-Thought:  i) CoT is effective in evaluating prompt following capability.
+Through CoT, the model more accurately discerns the significance of phrases such as “morphs into a mythical dragon”, allowing it to assign a more appropriate evaluation score.
+ii) CoT does not bring improvement in concept preservation evaluation.
+We argue that CoT may shift attention to unnecessarily important background or texture information, as shown in Fig.
 8.
-Figure 8: Case study of successful and failure case on DreamBench++ with Emu2.
 Table 6: Study of Chain-of-Though (CoT) and In-context Learning (ICL) on human alignment.
-Method TI [28] DreamBooth [86] DreamBooth-L [86] BLIP-D [55] Emu2 [95] IP-Adapt.-P [115] IP-Adapt. [115]
+Method TI DreamBooth DreamBooth-L BLIP-D Emu2 IP-Adapt.-P IP-Adapt.
 T2I Model SD v1.5 SD v1.5 SDXL v1.0 SD v1.5 SDXL v1.0 SDXL v1.0 SDXL v1.0
@@
 + 2 shot ICL -0.030 -0.002 -0.002 -0.051 -0.031 -0.155 -0.082
-4.3 Can we use free lunch to improve DreamBench++ evaluation?
-Table 6 shows the result of utilizing free lunch techniques, including chain-of-thought (CoT) [108] and In-Context Learning (ICL) [4, 12].
-CoT indicates that GPT-4 articulates its reasoning process before scoring, and ICL indicates GPT-4o is provided with human-written few-shot examples.
-Chain-of-Thought: i) CoT is effective in evaluating prompt following capability.
-Through CoT, the model more accurately discerns the significance of phrases such as “morphs into a mythical dragon”, allowing it to assign a more appropriate evaluation score.
-ii) CoT does not bring improvement in concept preservation evaluation.
-We argue that CoT may shift attention to unnecessarily important background or texture information, as shown in Fig.
-9.
+In-Context Learning:  ICL counterintuitively leads to a drop in alignment.
+This could be attributed to the patching scheme, sample selection, or inherent bias within GPT-4o, making it non-trivial to prompt effectively.
+Thus, we provide our detailed prompt and hope to inspire future works.
 Figure 8: Case study on CoT Prompting.
@@
 (b) However, it may also hinder visual concept preservation by drifting GPT’s attention away from the subject.
-In-Context Learning: ICL counterintuitively leads to a drop in alignment.
-This could be attributed to the patching scheme, sample selection, or inherent bias within GPT-4o, making it non-trivial to prompt effectively.
-Thus, we provide our detailed prompt and hope to inspire future works.
+4.4 How transferable is DreamBench++ across MLLMs?
+Table 7 reveals: i) DreamBench++ demonstrates robust generalization across different multimodal LLMs.
+The consistency ratios remain stable across various models, with Claude-3.5 achieving 74.00% and GPT-4o variants ranging from 67.22% to 84.23%.
+ii) The performance of MLLMs is continuously improving, as shown by the increasing Human-MLLM consistency in newer versions of GPT-4o (from 79.47% to 84.23%).
+This suggests that the accuracy of DreamBench++ evaluations may further improve as models advance.
+Table 7: Comparison of Human-MLLM consistency ratios. Ratio: the overall alignment between model evaluations and human evaluations, calculated as the average proportion of H-H consistency.
+Method TI DreamBooth DreamBooth-L BLIP-D Emu2 IP-Adapt.-P IP-Adapt. Ratio
+T2I Model SD v1.5 SD v1.5 SDXL v1.0 SD v1.5 SDXL v1.0 SDXL v1.0 SDXL v1.0
+H-H (Ground Truth) 0.685 0.647 0.656 0.613 0.746 0.602 0.591 100%
+GPT-4o-2024-05-13 (DreamBench++) 0.544 0.596 0.641 0.362 0.669 0.366 0.458 79.47%
+GPT-4o-2024-08-06 0.558 0.625 0.645 0.383 0.708 0.427 0.502 84.23%
+GPT-4o-mini-2024-07-18 0.496 0.575 0.538 0.274 0.702 0.238 0.289 67.22%
+Claude-3.5-Sonnet-20241022 0.500 0.583 0.623 0.354 0.625 0.278 0.427 74.00%
+Gemini-1.5-Pro-001 0.487 0.547 0.514 0.267 0.653 0.253 0.202 63.04%
 5 Related Works
-Personalized Image Generation aims to preserve concept consistency while accommodating the diverse contexts suggested by the instructions.
-In general, it can be traced back to early efforts on pixel-to-pixel (Pix2Pix) translation where the personalization orientation is free-form texts [11, 99, 75] or predefined translation across styles, seasons, species, or plants, etc [46, 127, 118, 106, 87].
+Personalized Image Generation
+aims to preserve concept consistency while accommodating the diverse contexts suggested by the instructions.
+In general, it can be traced back to early efforts on pixel-to-pixel (Pix2Pix) translation where the personalization orientation is free-form texts (Brooks et al., 2023; Tumanyan et al., 2023; Parmar et al., 2023) or predefined translation across styles, seasons, species, or plants, etc (Isola et al., 2017; Zhu et al., 2017; Zhang et al., 2023a; Wang et al., 2018; Saharia et al., 2022a).
 Modern efforts go beyond Pix2Pix translation toward a free-form image generation conditioned on both reference images and prompts.
-Some works focus on fine-tuning techniques that turn a general T2I model into a specialist personalization model [28, 86, 53, 92, 74] using LoRA [42] or contrastive learning [117, 35], learning the subject or style information by reconstructive autoencoding [102, 23].
+Some works focus on fine-tuning techniques that turn a general T2I model into a specialist personalization model (Gal et al., 2023a; Ruiz et al., 2023; Kumari et al., 2023; Sohn et al., 2023; Park et al., 2024) using LoRA (Hu et al., 2022) or contrastive learning (Zhang et al., 2022; He et al., 2020), learning the subject or style information by reconstructive autoencoding (Vincent et al., 2008; Dong et al., 2023).
 However, the necessity to fine-tune for new subjects limits their scalability.
 In contrast, encoder-based methods can generate subject-guided or style-guided images or edit images following prompts with one shot.
-Encoder- or adapter-based methods [125, 115, 109, 55, 48, 29, 16, 103, 105] train an encoder to encode the conditional image into embeddings, which are integrated into cross-attention mechanism in the diffusion process [40, 93, 68, 94].
-Adapter-free methods [66, 64, 37, 36, 11] extract the information, such as attention maps [36, RBModulation24] from reference images, and fuse them into the image generation process.
-Furthermore, multimodal large language models (MLLMs) that are trained on extensive multimodal sequences can also serve as general foundation models [24, 95, 73, 30].
-Benchmarking Image Generation involves a variety of metrics that focus on different aspects.
-Inception Score [89] and FID [39] judge image quality, while LPIPS [120], DreamSim [26], CLIP-I [81], and DINO Score [13] measure perceptual similarity.
-In text-guided generation, prompt-image alignment can be assessed by CLIP-T [81], CLIPScore [38], and BLIP Score [56, 57].
+Encoder- or adapter-based methods (Zheng et al., 2024; Ye et al., 2023; Wei et al., 2023; Li et al., 2023a; Jia et al., 2023; Gal et al., 2023b; Chen et al., 2023b; Wang et al., 2024a; b) train an encoder to encode the conditional image into embeddings, which are integrated into cross-attention mechanism in the diffusion process (Ho et al., 2020; Song et al., 2021a; Nichol & Dhariwal, 2021; Song et al., 2021b).
+Adapter-free methods (Lv et al., 2024; Liu et al., 2023d; Hertz et al., 2023b; a; Brooks et al., 2023) extract the information, such as attention maps (Hertz et al., 2023a) from reference images, and fuse them into the image generation process.
+Furthermore, multimodal large language models (MLLMs) that are trained on extensive multimodal sequences can also serve as general foundation models (Dong et al., 2024; Sun et al., 2024a; Pan et al., 2024; Ge et al., 2024).
+Additionally, some works (Wang et al., 2024b; Valevski et al., 2023; Yan et al., 2023; Ye et al., 2023; Xiao et al., 2023) focus on facial feature preservation.
+Benchmarking Image Generation
+involves a variety of metrics that focus on different aspects.
+Inception Score (Salimans et al., 2016) and FID (Heusel et al., 2017) judge image quality, while LPIPS (Zhang et al., 2018), DreamSim (Fu et al., 2023), CLIP-I (Radford et al., 2021), and DINO Score (Caron et al., 2021) measure perceptual similarity.
+In text-guided generation, prompt-image alignment can be assessed by CLIP-T (Radford et al., 2021), CLIPScore (Hessel et al., 2021), and BLIP Score (Li et al., 2022; 2023b).
 However, these metrics often fall short of reflecting human perception.
-To address this, human-aligned metrics [52, 112, 54] have been introduced, offering a more perceptive evaluation.
+To address this, human-aligned metrics (Ku et al., 2024; Xu et al., 2023; Lee et al., 2023) have been introduced, offering a more perceptive evaluation.
 Yet, they face limitations in scaling with the pace of new model developments.
-Thus, the necessity for automated and sustainable evaluation methods has emerged, with some [112, 60, 32] leveraging reward-model-based methods to encode human preferences, while others [51, 18, 110, 121, 43, 65] use multimodal  [12, 84, 5, 62] to automate the process and better mirror human tastes.
-While MLLM-based methods show promise in aligning with human preferences [121, 110, 45, 19], automated personalized evaluation remains an unresolved issue.
-VIEScore [51] assesses image generation quality by prompting GPT-4V [70] and LLaVA [62, 61], but is limited to four models in subject-driven tasks and obtains suboptimal results.
-Meanwhile, Dreambench [86], a common benchmark for personalized generative evaluation, only consists of 30 simple objects and lacks diversity comprehensiveness.
+Thus, the necessity for automated and sustainable evaluation methods has emerged, with some (Xu et al., 2023; Liang et al., 2023b; Guo et al., 2024) leveraging reward-model-based methods to encode human preferences, while others (Ku et al., 2023; Cho et al., 2023; Wu et al., 2024b; Zhang et al., 2023c; Hu et al., 2023; Lu et al., 2023) use multimodal  (Brown et al., 2020; Reid et al., 2024; Anil et al., 2023; Liu et al., 2023b) to automate the process and better mirror human tastes.
+While MLLM-based methods show promise in aligning with human preferences (Zhang et al., 2023c; Wu et al., 2024b; Huang et al., 2023; Cho et al., 2024), automated personalized evaluation remains an unresolved issue.
+VIEScore (Ku et al., 2023) assesses image generation quality by prompting GPT-4V (OpenAI, 2023) and LLaVA (Liu et al., 2023b; a), but is limited to four models in subject-driven tasks and obtains suboptimal results.
+Meanwhile, Dreambench (Ruiz et al., 2023), a common benchmark for personalized generative evaluation, only consists of 30 simple objects and lacks diversity comprehensiveness.
 6 Conclusions
 This paper introduces DreamBench++, a human-aligned personalized image generation benchmark.
-Extensive and comprehensive experiments have demonstrated our advantages in dataset diversity and complexity, as well as the alignment of our evaluation metrics with human preferences.
+Extensive and comprehensive experiments show significant advantages in dataset diversity and complexity, along with metrics that align with human preferences.
 In addition, we offer insights into prompt design for advanced multimodal GPTs, emphasizing the potential and challenges of enhancing GPT evaluation through chain-of-thought prompting and in-context learning.
 Our work aims to support future research on personalized image generation by providing a human-aligned benchmark and heuristics in utilizing advanced multimodal GPTs in visual evaluation.
-Appendix A Implementation Details
+Appendix A Qualitative Analysis
+With a more comprehensive and diverse collection of images, we have discovered numerous intriguing characteristics of these generation methods, as illustrated in Fig.
+9, that were not apparent on existing datasets such as DreamBench.
+Specifically, we observe that: i) Fine-tuning-based methods outperform encoder-based methods on images containing more subject-oriented information, such as an animal or object, as they preserve more intricate details in the generated images.
+However, for images containing a person, fine-tuning-based methods often fail to preserve facial and clothing features.
+This suggests that the personalized generation of human images is more demanding for visual concept preservation than for textual following capabilities, which is an advantage for encoder-based methods.
+ii) However, for style-oriented cases when subject details are less critical, encoder-based methods perform better than fine-tuning-based methods.
+This further highlights the strengths of encoder-based methods in that they are more adept at recognizing and extracting high-level visual semantics, including overall shape, style, and thematic features.
+Appendix B Implementation Details
 Table 8: Training hyperparameters on DreamBench and DreamBench++. BS: batch size, LR: learning rate, Steps: training steps.
@@
 Method T2I Model BS LR Steps BS LR Steps
-Textual Inversion [28] SD v1.5 4 5e-4 3000 1 5e-4 3000
-Dreambooth [86] SD v1.5 1 2.5e-6 1000 1 2.5e-6 250
-Dreambooth LoRA [86] SDXL v1.0 4 5e-5 500 1 5e-5 500
+Textual Inversion SD v1.5 4 5e-4 3000 1 5e-4 3000
+Dreambooth SD v1.5 1 2.5e-6 1000 1 2.5e-6 250
+Dreambooth LoRA SDXL v1.0 4 5e-5 500 1 5e-5 500
 The configurations for the training hyperparameters used in training-based methods on DreamBench and DreamBench++, are detailed in Table 8.
@@
 Method T2I Model Negative Prompt
-BLIP-Diffusion [55] SD v1.5 over-exposure, under-exposure, saturated, duplicate, out of frame, lowres, cropped,
+BLIP-Diffusion SD v1.5 over-exposure, under-exposure, saturated, duplicate, out of frame, lowres, cropped,
 worst quality, low quality, jpeg artifacts, morbid, mutilated, ugly, bad anatomy,
 bad proportions, deformed, blurry, duplicate
-IP-Adapter ViT-G [115] SDXL v1.0 deformed, ugly, wrong proportion, low res, bad anatomy, worst quality, low quality
-IP-Adapter-Plus ViT-H [115] SDXL v1.0 deformed, ugly, wrong proportion, low res, bad anatomy, worst quality, low quality
-We dedicate significant effort to tuning hyperparameters to ensure that the performance of each method on DreamBench is consistent with results reported in original papers.
-As shown in Table 9, the results of our reproduction are comparable to or even better than the official results reported by the original papers.
+IP-Adapter ViT-G SDXL v1.0 deformed, ugly, wrong proportion, low res, bad anatomy, worst quality, low quality
+IP-Adapter-Plus ViT-H SDXL v1.0 deformed, ugly, wrong proportion, low res, bad anatomy, worst quality, low quality
+We dedicate significant effort to tuning hyper-parameters to ensure that the performance of each method on DreamBench is consistent with results reported in original papers.
+Table 10 shows the results of our reproduction are comparable to or even better than the official results.
 Table 10: Reproduced results. Our reproduction is comparable to or better than the official results. N/A denotes that the official paper does not report the corresponding results.
@@
 […]
-Appendix B Limitation & Future Work
+Appendix C Additional Discussions
+C.1 Are multiple images for each instance necessary?
+In practice, multiple reference images are unnecessary for personalized image generation: i) The limited availability of reference images during daily usage makes single-image personalization more relevant.
+ii) Fine-tuning methods perform well with just one image, as shown in Appendix A.
+C.2 Which technical route performs better?
+Our comprehensive analysis reveals distinct advantages of different technical routes across various generation scenarios.
+For concept preservation, encoder-based methods demonstrate superior performance in human-centric generation, particularly in preserving precise facial features where fine-tuning approaches often struggle.
+This aligns with the recent trend in Identity-Preserving Generation tasks.
+However, fine-tuning methods show stronger capabilities in handling common objects and standardized appearances, while encoder-based approaches provide more consistent performance across diverse object types.
+The choice of base model significantly impacts generation quality.
+SDXL exhibits better concept preservation and prompt following capabilities compared to SD v1.5, primarily due to its more sophisticated text encoder architecture and diverse training data.
+Regarding prompt adherence, fine-tuning methods generally outperform encoder-based approaches by better preserving the original text-image conditional distribution through direct learning from image-text pairs.
+In contrast, encoder-based methods may face challenges in maintaining precise prompt alignment due to potential disruptions during feature injection.
+These findings suggest that the optimal technical route depends heavily on the specific application scenario.
+For applications prioritizing human identity preservation, encoder-based methods are preferable.
+However, for use cases requiring precise prompt following or handling common objects, fine-tuning approaches might be more suitable.
+This understanding helps guide future research directions in personalized image generation.
+Appendix D Limitation & Future Work
 Human-aligned evaluation & benchmarking is an emerging but challenging direction, and we have only made preliminary attempts at personalized image generation.
@@
 We believe that as visual world models continue to develop, the evaluation performance will be further optimized.
-Our future work will focus on more applications with human-aligned evaluation, such as 3D generation [78, 79, 63, 27], video generation [41, 91, 9], autonomous driving [33, 58, 119], and even embodied visual intelligence [59, 10, 25, 80, 123, 31].
+Our future work will focus on more applications with human-aligned evaluation, such as multimodal intelligence (Wei et al., 2024a; Zhu et al., 2025; Wei et al., 2024d; c; Zhu et al., 2024; Liu et al., 2024; Chen et al., 2024; Wei et al., 2024b; Gao et al., 2024a; b), audio generation (Gao et al., 2024c), 3D generation (Poole et al., 2023; Qi et al., 2023; Liu et al., 2023c; Gafni et al., 2022; Gao et al., 2023), video generation (Ho et al., 2022; Singer et al., 2023; Blattmann et al., 2023; Zhou et al., 2025), autonomous driving (Han et al., 2024; Li et al., 2023c; Zhang et al., 2023b), and even embodied visual intelligence (Goyal et al., 2022; Liang et al., 2023a; Brohan et al., 2023; Driess et al., 2023; Qi et al., 2025; Zhao et al., 2024; He et al., 2024b; Yu et al., 2024; 2025).
 Broader Impact